OpenStudy (anonymous):

Solve 5^(x + 5 )= 9x Solving Exponential Equations with Unequal Bases

OpenStudy (anonymous):

answer options: x ≈ 13.69 x ≈ 0.07 x ≈ –0.28 x ≈ –3.63

OpenStudy (amistre64):

the answers dont explain how ....

OpenStudy (amistre64):

youre best bet might be to get them to speak in natural logs and work it from that angle

OpenStudy (amistre64):

is that a 9x or a 9^x ?

OpenStudy (anonymous):

a 9^x sorry wrote it wrong.

OpenStudy (amistre64):

5^(x + 5 )= 9^x ; ln it all ln(5^(x + 5 ))= ln(9^x) ; lns pull out exponents (x+5) ln(5)= x ln(9) ; get the xs to one side and the rest to the other (x+5)/x = ln(9)/ln(5) ; assuming x not equal 0

OpenStudy (amistre64):

1+5/x = ln ... 5/x = ln .... 5/ln ... = x

OpenStudy (amistre64):

ugh, i forgot to type in the -1 :/ bad fingers bad!!

OpenStudy (amistre64):

1+5/x = ln ... 5/x = ln .... - 1 5/ln .. - 1 = x

OpenStudy (anonymous):

you need to take log of both sides and then try to solve for x using the properties of log